in his coin box, Brian has 12 fewer nickles than dimes. the value of his nickles and dimes is $2.40. how many each type does he have?
Answer by jorel1380(3719)

5n+10d=240
d=n+12
5n+10(n+12)=240
15n+120=240
15n=120
n=8
d=20
Brian has 8 nickels and 20 dimes..